OpenStudy (anonymous):

Find the measure of an angle if it's measure is 5 degrees less than 4 times the measure of it's complement.

OpenStudy (anonymous):

so (x+5) = 4*(90-x) (x+5)=4*(90-x) x+5 = 360-4x x=71

OpenStudy (anonymous):

Thanks!